Suppose the following equations represent an individual’s utility maximization problem: U(X,Y) = X0.5 + Y0.5 And the budget constraint is: I = PxX + PyY (a) Set up the individual’s maximization problem using the Lagrange technique. (b) Find the individual’s demand function for X and Y (Derive from first order condition). (c) Find the indirect utility function. (d) Find the expenditure function. (e) Find the share of X and Y on expenditure. (f) Find the marginal utility of income.
Answer #1

Date es tyo.s -0.5 olx ULX,Y)= xos I = Px X + Pyy h = xos tyos - x (Px X + Pu Y } -3) dL = 0.5X -XP x = 0 =) 0.5x10.se MPx -0

] - I = x P + P * 2. * (PxPy tra Yo Te pyta ta?) a taifa y

Date t ila Palute uLX,Y) = xot yos luf = | IPy jois- PapytPx 0.5 luf - (IPy) Pyos + (SPM) O spois Pris pois (Pvtly jois luf